OpenStudy (anonymous):

Find the equation in standard form which passes through (-2, -7) and (1, -1).

OpenStudy (binary3i):

m = {-7-(-1)}/{-2-1} = 2 y = mx +c y = 2x + c now put any point at the place of x and y -1 = 2 + c so c = -3 and equation is y = 2x -3